r=k*Ca*Cb. What are the units for the rate constant k?
Answer #1

For r- . TA EIB unit of r = MollL.5-1 TA = mol/L CBF mol/L to= r/TA.T13 - moll2 57 molt. 2.5t (mol/L) 2 unit of k= molt 2.51